OpenStudy (anonymous):

Short Answer: Simplify the following expression; express in terms of i. square root of -121

hartnn (hartnn):

whats square root of 121 ?

OpenStudy (anonymous):

11

hartnn (hartnn):

good., now \(\huge \sqrt{-a^2}=\sqrt{-1}\sqrt{a^2}=ia\) so what will be square root of -121 ?

hartnn (hartnn):

put a=11

OpenStudy (anonymous):

-11

hartnn (hartnn):

nopes, it should be 11i.

OpenStudy (anonymous):

so that is the answer 11i

hartnn (hartnn):

yup. did u understand how ?

OpenStudy (anonymous):

yeah you just take the square root and add i lol

hartnn (hartnn):

i added i because square root of -1 is i.

hartnn (hartnn):

and infact i multiplied 11 by i

OpenStudy (anonymous):

ok thank you